# Teleportation=Translation: Continuous recovery of black hole information

### v3

@@ 4.2 Geometric Derivation via Canonical Path
By definition (see Definition 3.9), the unitary path is $\tilde{U}(s):=J_{\tilde{\mathcal{M}}}\,J_{\tilde{\mathcal{N}}(s)}$. Its derivative at $s=0$ defines $\tilde{G}$ on the common core domain $\mathcal{D}_{\text{core}}$:
$\tilde{G}\;:=\;i\,\frac{d\tilde{U}(s)}{ds}\Big|_{s=0}\;=\;i\,J_{\tilde{\mathcal{M}}}\,\frac{d\,J_{\tilde{\mathcal{N}}(s)}}{ds}\Big|_{s=0}~,$ (30)
in accordance with (19). Here $J_{\tilde{\mathcal{M}}}$ is the modular conjugation for $\tilde{\mathcal{M}}$, and $J_{\tilde{\mathcal{N}}(s)}$ is that for the algebra $\tilde{\mathcal{N}}(s)$.
For a generic perturbation, the variation of $J$ usually involves complex spectral integrals of the modular operator arising from non-commutativity. However, a remarkable feature of the canonical $L^{p}$-interpolation path is that it follows the geodesic of the underlying modular geometry. Consequently, the complex non-commutative terms cancel out, [^2] Explicitly, the path is generated by the analytic scaling $\Delta_{s}^{it}=h^{ist}\Delta_{0}^{it}$. The variation with respect to $s$ extracts the generator $\ln h$ linearly, avoiding the parameter-integral convolution typical of generic perturbations. leaving a simple geometric proportionality:
$i\,J_{\tilde{\mathcal{M}}}\,\frac{dJ(s)}{ds}\Big|_{s=0}\;=\;-\,2\,\frac{dK(s)}{ds}\Big|_{s=0}~.$ (31)
Physically, this factor of 2 is a direct consequence of the geometric role of $J$. Since the modular conjugation implements a geometric reflection (effectively inverting the emergent spacetime), shifting the reflection axis along the geodesic results in a net displacement of twice that magnitude. This geometric logic is inherent to the canonical $L^{p}$ interpolation and justifies the specific coefficient in our operator identity.
Substituting the result of Lemma 4.1 ($K^{\prime}(0)=-P$) into (31), we obtain:
$\tilde{G}\;=\;-\,2\,K^{\prime}(0)\;=\;-\,2(-P)\;=\;2P~,$ (32)
establishing the algebraic identity on the dense domain $\mathcal{D}_{\text{core}}$. Crucially, this exact reconstruction of $2P$ holds regardless of the commutation relations between $K_{\tilde{\mathcal{M}}}$ and $P$, confirming that the result is a robust consequence of the modular structure rather than an artifact of specific symmetries.
It remains to lift the equality $\tilde{G}=2P$ from the core domain to the level of closed operators. By construction, the common analytic core $\mathcal{D}_{\text{core}}$ is invariant under the modular automorphism groups of both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Consequently, $\mathcal{D}_{\text{core}}$ consists of analytic vectors for the modular Hamiltonians $K_{\tilde{\mathcal{M}}}$ and $K_{\tilde{\mathcal{N}}}$, and thereby for their difference $P$. Thus, $2P$ is essentially self-adjoint on this domain. Similarly, $\tilde{G}$ is essentially self-adjoint on $\mathcal{D}_{\text{core}}$ by Nelson’s analytic vector theorem (Lemma 3.10). Since $\tilde{G}$ and $2P$ coincide on a common core where both are essentially self-adjoint, their unique self-adjoint closures are identical. This completes the proof. ∎
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A central physical question is whether the identity $\tilde{G}=2P$ represents a singular coincidence valid only at the boundary $s=0$, or a structurally robust property of the interpolation. Specifically, we must ensure that the generator does not exhibit large or uncontrolled fluctuations under small perturbations of the interpolation parameter $s$. We address this by employing Kosaki’s non-commutative $L^{p}$ theory Kosaki84 ; Kato , which provides the canonical geometric framework for analyzing perturbations in Type II von Neumann algebras.
@@
The generator $\tilde{G}$ exhibits structural stability in the strong resolvent sense. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. For any fixed complex number $z$ in the resolvent set $\rho(2P)$ (i.e., $z\notin\sigma(2P)$) and any state vector $\xi$ within the dense core domain $\mathcal{D}_{\text{core}}$, the deviation of the resolvent is quantitatively bounded by the variation of the associated projection $e_{\tilde{\mathcal{N}}(s)}$:
@@
Here, $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator, and $||\cdot||$ represents the vector norm on the Hilbert space (or the operator norm for the projection difference). The coefficient $C_{z}$ is a stability factor that scales inversely with the distance to the spectrum $\sigma(2P)$ (i.e., $C_{z}\sim\operatorname{dist}(z,\sigma(2P))^{-1}$).
@@
The proof establishes the local stability by applying perturbation theory for unbounded operators.
@@
  2. 2.
    Kato-Type Bound: We employ the standard first-order perturbation estimate for resolvents. The perturbation of $\tilde{G}(s)$ away from $\tilde{G}(0)=2P$ is effectively controlled by the distance between the projections. The constant $C_{z}$ is bounded by the inverse distance to the spectrum, $C_{z}\lesssim 1/|\operatorname{Im}(z)|$ Kato .
@@
    Quantification: For the canonical path, the projection perturbation scales linearly with $s$ near the origin: $e_{\tilde{\mathcal{N}}(s)}-e_{\tilde{\mathcal{M}}}||=O(s)$.
This confirms that the generator $\tilde{G}(s)$ converges to $2P$ in the strong resolvent sense as $s\to 0$, affirming the structural robustness of the $\tilde{G}=2P$ identity against small deformations. ∎
@@ 4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ posits a direct equivalence between the algebraic shift derived from modular theory and the geometric translation of the emergent spacetime. While operator-level identities are mathematically rigid, their physical validity in complex quantum gravity models requires verification through observables. We propose a concrete test using two-point correlation functions within the AdS/CFT correspondence.
We frame this test in the context of a two-sided eternal black hole, holographically dual to the thermofield double state $|\text{TFD}\rangle$. The local algebras $\mathcal{M}$ and $\mathcal{N}$ correspond to nested subregions of the boundary CFT, specifically associated with the Right exterior region. Let $O_{L}$ and $O_{R}$ be local probe operators acting on the Left and Right boundaries, respectively. Since the canonical shift $\tilde{U}(s)$ is generated by the algebra on the Right, it acts non-trivially on $O_{R}$.
@@
This correlator probes the “distance” through the wormhole between the fixed Left operator and the algebraically shifted Right operator. Since $\tilde{G}$ is defined as the infinitesimal generator of the path at $s=0$ (satisfying the expansion $\tilde{U}(s)=\mathbf{1}-is\tilde{G}+O(s^{2})$), the initial derivative determines the prediction for the operator identity. If $\tilde{G}=2P$ holds, we obtain:
@@
Eq. (35) provides the precise physical interpretation that the algebraic generator $\tilde{G}$ acting on the boundary operator induces a shift exactly twice that induced by the geometric translation generator $P$.
In the bulk gravity description, $P$ generates a geometric time translation along the horizon. Therefore, verifying Eq. (35) amounts to checking if the algebraic shift $\tilde{U}(s)$ effectively displaces the position of the horizon (or the stretched horizon) relative to the probe $O_{L}$ by a coordinate distance of $2s$.
This geometric action is deeply connected to the mechanics of holographic teleportation. In models like the Gao-Jafferis-Wall (GJW) protocol Gao2017 , the traversability is achieved by a gravitational shockwave that shifts the horizon. Here, our result implies that the algebraic generator $\tilde{G}$ explicitly implements this horizon shift without needing an external matter source, effectively simulating the backreaction required to bridge the entanglement wedge.
In semiclassical models like Jackiw-Teitelboim (JT) gravity Saad:2019lba , the quantity $2i\langle O_{L}[P,O_{R}]\rangle$ precisely quantifies this shift, manifesting as a Shapiro time delay for signals crossing the wormhole Shenker2014 ; Maldacena2016 . Thus, the factor of 2 confirms that $\tilde{U}(s)$ physically manipulates the geometry of the stretched horizon, facilitating the information transfer.
@@ 5 Conclusion
We have presented a framework connecting the discrete, algebraic teleportation protocol of vdH-V vdHV with the continuous, geometric picture provided by modular flow. In Type III settings, a key challenge is to construct a path that is both mathematically controlled and physically consistent. Our construction addresses this by defining a strongly continuous unitary path $\tilde{U}(s)$ with a well-defined self-adjoint generator $\tilde{G}$, satisfying the relation
@@
The unitary $\tilde{U}(s)\approx e^{-is\tilde{G}}$ describes an effective transport of quantum information. At $s=0$, the information resides in the relative commutant $\tilde{\mathcal{A}}=\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (corresponding to the island region just behind the horizon) and is dynamically inaccessible. As $s$ increases, the operator smoothly transports this information into the accessible radiation sector $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where $\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$. The endpoint corresponds to the canonical shift unitary $U_{\tilde{\Gamma}}$, where the factor of $2$ arises because $U_{\tilde{\Gamma}}$ is generated by two modular conjugations, effectively doubling the modular distance. In this view, the recovery of information via teleportation becomes algebraically analogous to a geometric spacetime translation generated by $P$, giving a precise form to the correspondence Teleportation=Translation.
@@
This perspective offers a consistent framework for approaching the black hole information paradox. Within this picture, global evolution is unitary: the operator $\tilde{U}(s)$ transports interior information to the exterior algebra rather than destroying it, suggesting that no information is lost during the evaporation process. No-cloning is naturally preserved by algebraic complementarity, as the information is encoded in distinct algebras across different values of $s$ without duplication. The apparent thermality can be interpreted as a consequence of the Type III algebraic structure—specifically the absence of a tracial state—rather than an intrinsic feature of the global dynamics. Furthermore, lifting the description to the semifinite envelope clarifies that entropy and thermality are governed by modular, as opposed to tracial, structures.
The property of dynamic idempotency induces a nested flow of algebras, offering a controlled, continuous realization of the Page/Hayden–Preskill decoding protocol. This framework shares key conceptual features with holographic quantum error correction and entanglement wedge reconstruction. In particular, the smooth transport of information from the initially inaccessible relative commutant to the radiation sector is analogous to the recoverability conditions in error-correcting codes Almheiri_2015 . Similarly, the generator $\tilde{G}=2P$ provides an algebraic description of accessing the island interior, paralleling the geometric reconstruction of the entanglement wedge Penington19 .
[…]
In summary, this formal development indicates that continuous, idempotent, and modular-covariant teleportation is algebraically equivalent to a spacetime translation. This identification provides a unitary and calculable channel for information transfer, supporting the view that black holes may evaporate in accordance with the principles of quantum mechanics and algebraic locality.
@@
It is worth noting that these results are derived within an algebraic QFT framework on a fixed background spacetime. While we demonstrate that information flow remains unitary and consistent with geometric translation in this setting, this work primarily provides a kinematic framework for information recovery; the full quantum gravitational dynamics, including the back-reaction of the evaporation process on the geometry, remains an open and important frontier.
@@
Our work suggests several directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Since $P$ generates translations across the horizon, our continuous protocol may provide an algebraic description of the GJW protocol Gao2017 , potentially offering a non-perturbative perspective on the wormhole opening mechanism via modular flow. Second, verifying the correlation function equality $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or JT gravity, would help bridge our algebraic results with semiclassical gravity calculations.
@@
Finally, extending this framework to incorporate gravitational back-reaction is a critical next step. Since our protocol $\tilde{U}(s)$ is explicitly constructed within the crossed-product envelope—a structure known to capture $1/N$ corrections and observer energy constraints Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET —it offers a natural starting point for such an extension. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a dynamic theory of emergent spacetime geometry from quantum entanglement.

### v4

@@ 4.2 Geometric Derivation via Canonical Path
We derive the generator explicitly by utilizing the Type II nature of the crossed product algebra $\tilde{\mathcal{M}}$. Recall from Definition 3.8 that the unitary path is defined as $\tilde{U}(s)=J(0)J(s)$, where $J(0):=J_{\tilde{\mathcal{M}}}$ is the fixed reference modular conjugation, and $J(s):=J_{\tilde{\mathcal{N}}(s)}$ is the conjugation associated with the interpolating subalgebra $\tilde{\mathcal{N}}(s)$. The generator is given by
$\tilde{G}=iJ(0)\frac{dJ(s)}{ds}\Big|_{s=0}~.$ (30)
In the crossed product construction, the dual weights $\tilde{\phi}_{s}$ act as weights on a Type II von Neumann algebra. Consequently, they can be represented via the canonical trace $\tau$ and the Radon-Nikodym derivative $h$. Specifically, the density operator for the weight $\tilde{\phi}_{s}$ is $h^{s}$ (relative to the reference weight $\tilde{\phi}_{0}$ corresponding to $h^{0}=\mathbf{1}$). According to the standard transformation rules for modular conjugations in Type II settings, the conjugation $J(s)$ associated with the weight $\tilde{\phi}_{s}(\cdot)=\tau(h^{s}\,\cdot\,)$ is related to the reference conjugation $J(0)$ by the unitary transformation generated by the density operator
$J(s)=h^{is}J(0)h^{-is}~.$ (31)
From Lemma 4.1, we identified the logarithm of the Radon-Nikodym derivative as the modular momentum: $\ln h=P$. Substituting this into (31), we obtain the explicit time-evolution of the modular conjugation
$J(s)=e^{isP}J(0)e^{-isP}~.$ (32)
We now differentiate this expression at $s=0$ to find $J^{\prime}(0)$:
$J^{\prime}(0)$ $=\frac{d}{ds}\left(e^{isP}J(0)e^{-isP}\right)\Big|_{s=0}$
$=(iP)J(0)+J(0)(-iP)~.$ (33)
Substituting $J^{\prime}(0)$ back into the definition of $\tilde{G}$ in (30) yields
$\tilde{G}=-J(0)PJ(0)+J(0)^{2}P~.$ (34)
This expression simplifies immediately by invoking the structural properties of the modular conjugation. Specifically, $J(0)$ is an involution ($J(0)^{2}=\mathbf{1}$) and acts as an anti-unitary reflection on the modular Hamiltonian, implying the CPT inversion symmetry J(0)PJ(0)=PJ(0)PJ(0)=-P. Applying these identities, we obtain
$\tilde{G}=-(-P)+\mathbf{1}\cdot P=2P~.$ (35)
This algebraic derivation confirms the identity $\tilde{G}=2P$ on the common core $\mathcal{D}_{\text{core}}$. Since $\mathcal{D}_{\text{core}}$ is a core for both $\tilde{G}$ (by Nelson’s theorem) and $P$ (by modular covariance), the identity holds for their unique self-adjoint closures. ∎
@@ 4.3 Stability Quantification via Non-Commutative $L^{p}$ Spaces
A central physical question is whether the identity $\tilde{G}=2P$ represents a singular coincidence valid only at the boundary $s=0$, or a structurally robust property of the interpolation. Specifically, we must ensure that the generator does not exhibit large or uncontrolled fluctuations under small perturbations of the interpolation parameter $s$. We address this by employing the theory of non-commutative $L^{p}$ spaces Kosaki84 ; PisierXu , which provides the canonical geometric framework for analyzing perturbations in the Type II crossed product algebra.
@@
The generator $\tilde{G}$ exhibits structural stability in the Kato-type strong resolvent sense. Let $\tilde{G}(s)$ denote the instantaneous generator of the unitary path at parameter $s$. For any fixed complex number $z$ in the resolvent set $\rho(2P)$ (defined as the complement of the spectrum $\sigma(2P)$ in the complex plane) and any state vector $\xi$ within the dense core domain $\mathcal{D}_{\text{core}}$, the deviation of the resolvent from the modular momentum limit is linearly bounded by the perturbative interpolating parameter $s$:
@@
Here, $R_{T}(z)=(T-z)^{-1}$ denotes the resolvent operator, and $||\cdot||$ represents the vector norm on the Hilbert space. The coefficient $C_{z}$ is a stability factor that scales inversely with the distance to the spectrum $\sigma(2P)$ (i.e., $C_{z}\sim\operatorname{dist}(z,\sigma(2P))^{-1}$), indicating that stability is most sensitive near the spectral values. This inequality implies that $\tilde{G}(s)$ converges to $2P$ in the strong resolvent sense as $s\to 0$. Crucially, by the Trotter-Kato theorem, this guarantees the convergence of the generated unitary groups, ensuring that the physical dynamics of $\tilde{G}(s)$ smoothly approach those of $2P$ (strong convergence of dynamics).
@@
The proof relies on the analytic rigidity of the canonical interpolation path within the Type II setting, utilizing the specific geometry of non-commutative $L^{p}$ spaces.
@@
    Geometric Foundation (Uniform Convexity $\to$ Analyticity): The interpolation path is constructed within non-commutative $L^{p}$ spaces ($1<p<\infty$) associated with the crossed product algebra. A crucial feature of these spaces is their uniform convexity Kosaki84 . Geometrically, uniform convexity implies that the unit ball is strictly rotund without flat regions. In the context of complex interpolation, this geometric rigidity ensures that the map $s\mapsto\tilde{U}(s)$ is not merely continuous but real-analytic for $s\in(0,1)$ and differentiable at the boundary $s=0$.
  2. 2.
    From Differentiability to Linear Bound: The analyticity established above guarantees that the generator $\tilde{G}(s)$ is a differentiable function of $s$ at the origin. By the mean value theorem for operator-valued functions, the deviation of the generator from its limit is bounded linearly by the parameter $s$:
@@
    for some constant $C$ related to the derivative $\tilde{G}^{\prime}(0)$. This linear bound ($\mathcal{O}(|s|)$) is a direct consequence of the smoothness of the geometry; without uniform convexity, the path could be fractal or non-differentiable, violating this bound.
@@
    Resolvent Estimate: To translate the generator bound to the resolvent bound, we employ the second resolvent identity:
    $R_{\tilde{G}(s)}(z)-R_{2P}(z)=R_{\tilde{G}(s)}(z)\big(2P-\tilde{G}(s)\big)R_{2P}(z)~.$ (38)
    Taking the norm and applying the linear bound from the previous step yields:
[…]
    where $C_{z}^{\prime}$ represents the uniform bound on the operator norm of the perturbed resolvent $R_{\tilde{G}(s)}(z)||_{\textrm{op}}$ for small $s$. The composite stability factor $C_{z}$ encapsulates the norms of the resolvents (scaling as $\text{dist}(z,\sigma)^{-1}$) along with the geometric derivative constant $C$.
This derivation confirms that the structural robustness ($\tilde{G}=2P$) is not accidental but grounded in the uniform convexity of the underlying modular geometry, which enforces the linear control of fluctuations. ∎
@@ 4.4 Correlation-Function Test of the Conjecture
The operator identity $\tilde{G}=2P$ posits a direct equivalence between the algebraic shift derived from modular theory and the geometric translation of the emergent spacetime. While operator-level identities are mathematically rigid, their physical validity in complex quantum gravity models requires verification through observables. We propose a specific test using two-point correlation functions within the framework of the AdS/CFT correspondence.
We frame this test in the context of a two-sided eternal black hole, holographically dual to the thermofield double state $|\text{TFD}\rangle$. To precisely model the information transport, we adopt the algebraic setup relevant to the black hole information paradox. We identify $\tilde{\mathcal{N}}$ with the algebra of the old black hole and $\tilde{\mathcal{M}}$ with the enlarged algebra that includes the entire black hole plus an infalling message (Alice’s diary). In this construction, the relative commutant $\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ represents the specific information of Alice’s diary. Technically, the canonical shift operator $\tilde{U}(s)$ is constructed to transport this information from $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the dual relative commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ (the radiation sector). According to our main theorem ($\tilde{G}=2P$), the infinitesimal generator responsible for moving Alice’s diary into the radiation sector coincides exactly with twice the generator of the geometric modular flow (the boost).
Let $O_{L}$ and $O_{R}$ be local probe operators acting on the Left and Right asymptotic boundaries, respectively, where $O_{R}$ is associated with the infalling message. Since the generator $\tilde{G}$ acts as a geometric boost generator on the Right wedge, the corresponding unitary $\tilde{U}(s)$ induces a non-trivial coordinate shift on the probe $O_{R}$ for small $s$, effectively displacing its position relative to the horizon, while leaving the Left operator $O_{L}$ invariant.
@@
This correlator probes the geodesic distance through the wormhole between the fixed Left operator and the algebraically shifted Right operator. Since $\tilde{G}$ is defined as the infinitesimal generator of the path at $s=0$ (satisfying the expansion $\tilde{U}(s)=\mathbf{1}-is\tilde{G}+O(s^{2})$), the initial response determines the prediction for the operator identity. If $\tilde{G}=2P$ holds, we obtain
@@
Eq. (41) provides a precise physical interpretation: the algebraic generator $\tilde{G}$, originally defined to extract Alice’s diary from the black hole, effectively induces a translation on the boundary operator $O_{R}$ with exactly twice the magnitude of the geometric generator $P$.
In the bulk gravity description, $P$ acts as the generator of a geometric boost. Crucially, this boost shifts the null coordinate $v$, effectively displacing the horizon inwards relative to the probe. This implies that a region previously hidden in the interior becomes accessible from the exterior. Therefore, verifying Eq. (41) involves checking whether the algebraic shift $\tilde{U}$ displaces the coordinate position of the message $O_{R}$ inward across the original horizon location by a coordinate distance twice the corresponding geometric translation.
This geometric action is deeply connected to the mechanics of holographic teleportation. In models like the Gao-Jafferis-Wall (GJW) protocol Gao2017 , traversability is achieved by a shockwave that shifts the horizon. Our result implies that the algebraic generator $\tilde{G}$ mimics this horizon shift intrinsically. Rather than requiring an external matter source, the algebraic flow $\tilde{U}(s)$ induces a frame transformation equivalent to the geometric displacement required to bridge the entanglement wedge and recover the information. In semiclassical models like Jackiw-Teitelboim (JT) gravity Saad:2019lba , the quantity $i\langle O_{L}[P,O_{R}]\rangle$ relates to the Shapiro time delay for signals crossing the wormhole Shenker2014 ; Maldacena2016 . Thus, the factor of 2 confirms that $\tilde{U}(s)$ physically manipulates the geometry of the stretched horizon, facilitating the information transfer.
@@ 5 Conclusion
In this work, we have presented a continuous canonical path connecting the discrete, algebraic teleportation protocol of vdH-V vdHV within the framework of algebraic quantum field theory characterized by Type III von Neumann algebras. A defining feature of Type III algebras is the absence of a normal, faithful, tracial state, which precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. To overcome this structural obstacle, we utilized the crossed product construction to lift the Type III inclusion $\mathcal{N}\subset\mathcal{M}$ to the semifinite enveloping algebras $\tilde{\mathcal{N}}\subset\tilde{\mathcal{M}}$.
Leveraging Haagerup-Kosaki non-commutative $L^{p}$-space theory, we have constructed the unitary canonical shift operator $\tilde{U}(s)$ for $s\in[0,1]$. This continuous unitary path describes the interpolated transport of quantum information: as $s$ increases, the operator smoothly transfers the information residing in the relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ into the accessible radiation sector $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ within the extension $\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$.
Our construction yields a strongly continuous unitary path $\tilde{U}(s)$ generated by a well-defined self-adjoint operator $\tilde{G}$, satisfying the relation
@@
Crucially, the factor of $2$ in this equation arises because the generator combines two modular conjugations; consequently, the effective displacement of the information corresponds to twice the geometric distance associated with a single modular reflection. In this view, information recovery via teleportation becomes algebraically equivalent to a geometric spacetime translation generated by $P$, providing a precise formulation of the correspondence Teleportation = Translation.
@@
This perspective offers a consistent framework for approaching the black hole information paradox. Within this picture, the recovery process is unitary: the operator $\tilde{U}(s)$ transports interior information to the exterior algebra rather than destroying it, suggesting that no information is lost during the evaporation process. No-cloning is naturally preserved by algebraic complementarity, as the information falling into the black hole is rapidly scrambled and subsequently teleported into the accessible radiation sector without duplication. The apparent thermality of black hole radiation is interpreted as a consequence of the Type III algebraic structure—specifically the absence of a tracial state—rather than an intrinsic feature of the global dynamics. This structural feature is directly tied to the presence of the horizon, which enforces the locality of QFT. Furthermore, lifting the description to the semifinite envelope unveils the algebraic structure underlying the Type III algebra, thereby clarifying the modular flow of information via the tracial structure of the extended Type II system.
@@
It is important to note that these results are derived within an algebraic QFT framework on a fixed background spacetime. This implies that the essential mechanism for resolving the black hole information paradox is accessible within the semiclassical approximation. We have demonstrated that, even in this semiclassical regime, the information flow remains unitary and consistent with geometric translation. On the other hand, the full quantum gravitational dynamics—particularly the back-reaction of the evaporation process on the geometry and the final stage of black hole evaporation—remain a critical frontier.
Our work suggests several directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Unlike the GJW protocol Gao2017 , which requires an external source (such as a shockwave) to shift the horizon, our continuous protocol generates a horizon shift intrinsically through modular interpolation. However, since both mechanisms rely on the generator $P$ to translate information across the horizon, incorporating additional algebraic conditions—such as specific coupling deformations—into our framework may provide a purely algebraic description of the GJW protocol. This would offer a non-perturbative perspective on the wormhole opening mechanism via modular flow.
@@
Second, verifying the correlation function relation $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or JT gravity, would help bridge our algebraic results with semiclassical gravity calculations.
@@
Finally, extending this framework to incorporate gravitational back-reaction represents a critical next step. Since our protocol $\tilde{U}(s)$ is explicitly constructed within the crossed-product envelope—a structure known to capture $1/N$ corrections and observer energy constraints Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET —it offers a natural starting point for such an extension. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a dynamical theory of emergent spacetime geometry arising from quantum entanglement.